Interactive Solution 22.61 offers one approach to problems such as this one. The secondary coil of a step-up transformer provides the voltage that operates an electrostatic air filter. The turns ratio of the transformer is 39:1. The primary coil is plugged into a standard 120-V outlet. The current in the secondary coil is 1.9 x 10-3 A. Find the power consumed by the air filter.
Answer #1

Given that -

Primary voltage Vp = 120 V

Secondary current Is = 1.9 x 10^-3 A

Again we know that Vp/Vs = Np/Ns

=> 120/Vs = 1/39 [turns ratio given is 39:1]

=> Vs = 39*120 = 4680 V

therefore power consumed by the air filter = Ps = 4680 V * 1.9 x 10^-3 A = 8.892 W